Circular No. 8792

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 A1 (LOVAS) Below are orbital elements by B. G. Marsden linking 38 observations

of P/1986 W1 = P/2007 A1 (cf. IAUC 8791; mean residual 0′′.7):

Epoch = 1986 Sept. 7.0 TT

T = 1986 Sept. 2.2572 TT  ω = 71.^{o}4102
e = 0.592752  Ω = 283.7467 }2000.0
q = 1.458819 AU  i = 1.5296

a = 3.582141 AU n^{o} = 0.1453751 P = 6.78 years Epoch = 2006 Dec. 11.0 TT

T = 2006 Dec. 12.4538 TT  ω = 77.^{o}4035
e = 0.604101  Ω = 278.3039 }2000.0
q = 1.395368 AU  i = 1.5532

a = 3.524554 AU n^{o} = 0.1489525 P = 6.62 years

SUPERNOVAE 2007A, 2007B, 2007C Three apparent supernovae have been discovered on unfiltered CCD

frames: 2007A by T. Puckett and T. Orff (cf. IAUC 8781; 0.35-m reflector at Ellijay, GA; source of tabulated data below) and by D. Madison and W. Li (LOSS/KAIT; cf. IAUC 8789); and 2007B (0.6-m reflector) and 2007C (0.3-m reflector) by K. Itagaki (via S. Nakano; cf. IAUC 8782).

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007A Jan. 2.02 0^{h}25^{m}16.^{s}66 +12^{o}53′12′′.5 16.0 1′′.2 W, 10′′.1 N
2007B Jan. 5.38 22 35 31.10 +34 48 06.6 16.7 7′′ W, 6′′ S
2007C Jan. 7.86 13 08 49.30 − 6 47 01.0 15.9 9′′ E, 22′′ S

Additional magnitudes for 2007A in NGC 105: 2005 Dec. 15, [19.1 (Puckett); 2006 Dec. 7.15, [19.5 (KAIT); 20.13, [18.3 (KAIT); 2007 Jan. 3.12, 16.3 (KAIT); 4.14, 16.0 (J. Newton, Portal, AZ, 0.40-m reflector). The LOSS team reports position end figures 16^{s}.68, 12′′.7 for 2007A. Additional magnitudes for 2007B in NGC 7315: 2006 Dec. 25.41, [19.0; 2007 Jan. 7.395, 16.4. Additional magnitudes for 2007C in NGC 4981: 2006 Dec. 23.87, [18.5 (Itagaki); 2007 Jan. 8.704, 16.0 (K. Kadota, Ageo, Japan, 0.25-cm f/5 reflector; position end figures 49^{s}.32, 47′00′′.3); 8.71, 15.8 (Itagaki). All three objects were found near maximum: 2007A and 2007B being type-Ia supernovae, and 2007C of type Ib (details on CBETs 796, 799, and 800).

2007 January 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green